OpenStudy (ashy98):

1.) Evaluate the following expression using the values given: Find x3 - 2y2 - 3x3 + z4 if x = 3, y = 5, and z = -3 A.) -185 D.) -77 C.) -23 D.) 85

OpenStudy (cj49):

27-50-243+243

OpenStudy (cj49):

-23

OpenStudy (ashy98):

i did that wrong my bad!

OpenStudy (cj49):

nopes i had made a mistake...-23 is the ans.

OpenStudy (ashy98):

okay thank you !(:

OpenStudy (anonymous):

Please try to use the equation editor when you have exponents. It is easier to read and easier to understand. \[X ^{3} - 2Y ^{2} - 3X ^{3} + Z ^{4}\]You could, if you wish, start by simplifying: Xcubed - 3 Xcubed is equal to minus 2 Xcubed but it's probably easier to just evaluate each term. each exponent only goes with the variable immediately in front of it (unless you have parentheses involved which we don't here) Look at \[- 2Y ^{2}\]How much is Y squared? Y is 5, so 5 squared is 25. Change Ysquared to 25 \[X ^{3} - 2(25) - 3X ^{3} + Z ^{4}\] Look at 3Xcubed? X is 3, so Xcubed is X*X*X = 3*3*3 = 27. Change Xcubed to 27 \[X ^{3} - 2(25) - 3(27) + Z ^{4}\] Can you finish it from here? Remember, when you calculate Z to the fourth power, that a negative number raised to an even power has an even result. -3 squared is equal to 9 because (-3)(-3) has two minuses, and in math, two wrongs always make a right. (When multiplying, two minus signs become a positive) So, -3 to the fourth power is (-3)(-3)(-3)(-3) = + what? I bet you can finish the problem.